Session: Feb 2006 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Torun (Kaz Borkowski at Fri Feb 24 12:37:22 2006) Minor failures : ANT
After repair of the telescope elevation drive we started this project normally, however at 3:50 UTC there was again an outage and the same problem arose with the drive. Observations were resumed (on a new DP, because of a problem with Mk5) at 8:50 only to meet with yet another crash at 11:35 because of a short power failure. Also, a few times the schedule did not allow enough time for the telescope to move to new source (this one concers other telescopes as well).

Onsala (Michael Lindqvist at Fri Feb 24 12:44:23 2006) Minor failures :
Lost a several scans since the source had to low elevation. When we restarted the FS after EF012 ended we got the following error 'ERROR sc -13 setcl: formatter to FS time difference 0.5 seconds or greater'.

Medicina (Fiocchi at Fri Feb 24 14:27:29 2006) Success :
Rain for all the experiment

Westerbork (Tony Foley at Mon Feb 27 12:49:25 2006) Minor failures :
Single dish was used. Please change the coordinates in the vexfile accordingly. Looks OK (but see note about single dish 'tracking' data being wrong up to 20seconds after a source change.

Effelsberg (A. Polatidis at Tue Feb 28 09:35:26 2006) Minor failures : RECE WIND
Stopped 0310-0450 UT due to snow. Receiver problem during the first half of the experiment, stopped 0730-1025 UT to investigate and fix. Source at very low elevation.

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:13:10 2006) Failure : RECO
The MkII telescope was used (Jb2). Offsets were incorrectly set hence sources probably out of telescope beam. Disk pack JOD-0023 subsequently developed "directory flawed or missing" error (during experiment ea035a). No data visible on disk pack but may be recoverable. Some brief elevation drive failures at low elevation.

Cambridge (Alastair Gunn at Fri Mar 3 17:13:54 2006) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 7 15:52:47 2006) Success :

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)